# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ XII Non-Equilibrium Quantum Sensing: Quantum Many-Body Scars and Sensing
are the scar states of Hamiltonian Eq. (77) with $\phi{=}\pm\pi/2$ Mark and Motrunich (2020). In other words, the eigenstates $|\psi(s)\rangle$, defined by $H(\phi{=}{\pm}\pi/2)|\psi(s)\rangle{=}h(s{-}N/2)|\psi(s)\rangle$, violate the eigenstate thermalization hypothesis as they show a subvolume-law entanglement growth. Rewriting the initial product state $|\psi_{t=0}\rangle{=}\otimes_{i=1}^{N}(|{+1}\rangle{+}|{-1}\rangle){/}\sqrt{2}$ in terms of the elements of the Dicke-scar subspace results in $|\psi_{t{=}0}\rangle{=}\sum_{s=0}^{N}\sqrt{\dfrac{N!}{2^{N}s!(N-s)!}}|\psi(s)\rangle$. The time-evolved state in this subspace is obtained as $|\psi_{t}^{*}\rangle{=}\otimes_{i=1}^{N}(e^{-it^{*}h/2}|{+1}\rangle+e^{it^{*}h/2}|{-1}\rangle){/}\sqrt{2}$ results in $\delta\lambda{=}1/\sqrt{Nt^{*}T}$, despite the presence of strong interaction. Therefore, the initial state which is prepared in the Dicke-scar subspace of Hamiltonian $H(\phi{=}\pm\pi/2)$ can result in a vanishing estimation error via its long-lived coherence property $t^{*}{\rightarrow}{\infty}$. More discussion on the advanced role of quantum many-body scar in quantum-enhanced sensing can be found in Refs. Dooley (2021); Dooley et al. (2023); Desaules et al. (2022).
@@ XIII Non-Equilibrium Quantum Sensing: Sequential Measurements Metrology
Measurement represents another distinct feature that sets apart the classical from the quantum world. Indeed, the concept of wave-function collapse due to projective measurements on the quantum system has sparked several discussions within quantum theory itself Machida and Namiki (1980a, b); Furry (1936); Zeh (1970). In quantum many-body probes with partial accessibility, measurements can only be conducted locally on a subsystem. However, in these scenarios, even though the measurement is local, the wave function collapse is global and affects the entire system’s wave function. This intriguing phenomenon has been the subject of intensive studies Busch et al. (1990); Schmidt and Gemmer (2020); Ban (2021); Skinner et al. (2019b); Block et al. (2022a); Benoist et al. (2023); Haapasalo et al. (2016); Benoist et al. (2019); Burgarth et al. (2014); Rybár and Ziman (2015), such as a unique feature for quenching many-body systems Bayat (2017); Pouyandeh et al. (2014); Ma et al. (2018); Bayat et al. (2018), leading to potential new types of phase transitions Choi et al. (2020); Feng et al. (2023); Minato et al. (2022); Skinner et al. (2019a); Gullans and Huse (2020); Block et al. (2022b). These measurement-induced phase transitions have also been recently linked with quantum enhanced sensitivity (Paviglianiti and Silva, 2023; Di Fresco et al., 2024).
In conventional sensing strategies, data is collected through independent and identical (IID) probability distributions. This means that after measuring the probe, one needs to reset it to its original quantum state for another round of experiments or, equivalently, use identical copies. Note that the formulation of the Cramér-Rao inequality (see Eq. (2)) assumes the resetting of the probe after each measurement or equivalently using $M$ identical probes at once. A simple yet versatile sensing strategy has been devised using sequential measurements on the probe at regular time intervals Burgarth et al. (2015); Guta and Kiukas (2017); Catana et al. (2015); Montenegro et al. (2022); Mabuchi (1996a); van Horssen and Guţă (2015). This innovative sequential sensing scheme is of significant importance in quantum metrology for several reasons: (i) it avoids extremely time-consuming resetting times, (ii) it avoids the need for highly correlated or complex initial states, and (iii) available experimental measurements suffice for evidencing quantum-enhanced sensitivity. The sequential measurement sensing scheme, see Fig. 17, is an iterative sensing protocol:
@@
    At time $\tau_{i}$ a positive operator-valued measure (POVM) $\Upsilon_{\gamma_{i}}=\Pi^{\dagger}_{\gamma_{i}}\Pi_{\gamma_{i}}$, where $\Upsilon_{\gamma_{i}}\geq 0$ and $\{\Upsilon_{\gamma_{i}}\}$ are the elements of the POVM with random outcome $\gamma_{i}$, is performed on the probe, updating the quantum state into
    $\rho^{(i{+}1)}(0){=}\frac{\Pi_{\gamma_{i}}\rho^{(i)}(\tau_{i})\Pi_{\gamma_{i}}^{\dagger}}{p(\gamma_{i})},$ (80)
@@
    $p(\gamma_{i}){=}\text{Tr}[\Pi_{\gamma_{i}}\rho^{(i)}(\tau_{i})\Pi_{\gamma_{i}}^{\dagger}],$ (81)
@@
In general, the data collected from sequential measurements are non-IID, meaning that the measurement data are not independent of each other—i.e., one deals with effectively different probes at each measurement step. It is essential to emphasize that, unlike standard sensing schemes, there is no need for specific maximally entangled probe, the need for supporting quantum phases, feedback mechanisms, or quantum control. Naturally, standard sensing scheme reduces to the particular case of $n_{\mathrm{seq}}{=}1$.
Figure 17: Comparison between standard and sequential measurements sensing strategies: $(a)$ Standard strategy for estimating a parameter $g$ of a quantum system, where measurement data $\{s_{1},\ldots,s_{N}\}$ are collected by independent and identical experiments. Every time the experiment is performed, the system is reset to some specific known initial state $\rho_{0}$. $(b)$ Sequential scheme for estimating a parameter $g$ of a quantum system, where the measurements are performed sequentially to collect data $\{s_{1},\ldots,s_{N}\}$ without resetting the state of the system every after the measurement and the initial state $\rho_{0}$ can be arbitrary. Figure taken from Burgarth et al. (2015).
By harnessing the sequential measurements followed by free evolution at regular time intervals, the sequential sensing scheme has led to significant advancements, including Hamiltonian identification Mabuchi (1996b), sequential measurements sensing schemes Li and Cui (2023); O’Connor et al. (2024); Clark et al. (2019a); De Pasquale et al. (2017a); Clark et al. (2019b); Rasbi et al. (2022); Burgarth et al. (2015); Catana et al. (2015); Guţă (2011); De Pasquale et al. (2017b); Ritboon et al. (2022); Bompais and Amini (2023); Bompais et al. (2022); Gherardini et al. (2019); Gambetta and Wiseman (2001a); Müller et al. (2016); Nagali et al. (2012); Kiilerich and Mølmer (2015).
@@
$H=-J\sum_{j=1}^{N-1}\bm{\sigma}_{j}\cdot\bm{\sigma}_{j+1}+B_{x}\sigma_{1}^{x}+B_{z}\sigma_{1}^{z},$ (82)
where $\bm{\sigma}_{j}{=}(\sigma_{j}^{x}{,}\sigma_{j}^{y}{,}\sigma_{j}^{z})$ represents a vector composed of Pauli matrices acting on qubit site $j$, $J$ denotes the exchange interaction, and $\bm{B}{=}(B_{x}{,}0{,}B_{z})$ stands for the local magnetic field to be estimated located at first site. For the sake of simplicity, a ferromagnetic state $|\psi(0)\rangle=|\downarrow\downarrow\downarrow\ldots\rangle$ is probed at site $N$ in the $\sigma_{z}$ basis at regular time intervals $J\tau=N$. The latter allows enough time for the quantum correlations to spread throughout the spin chain. It has been shown Montenegro et al. (2022) that a synchronization-like magnetization dynamics occurs between the first and last sites. This opens up the possibility of extracting information about the local magnetic field (located at the first site) by observing the dynamics of the readout qubit (located at the last $N$ site). Remote sensing holds special significance in biology, where non-invasive sampling of organic tissues is desirable to protect the biological material (see quantum illumination techniques Albarelli et al. (2020); Casariego et al. (2022)).
@@
$T=M(t_{\mathrm{init}}+t_{\mathrm{evo}}+t_{\mathrm{meas}}n_{\mathrm{seq}}),$ (83)
@@
Despite the above, the investigation of sequential measurement sensing schemes is usually limited to short lengths of measurement sequences due to the exponential growth of measurement outcomes with the number of sequences—$(2^{N})^{q}$ for $N$ measurements on $q$ spins. Thus, evaluating all probabilities distributions becomes intractable. For estimating sensing precision with a large number of sequential measurements, indirect approaches have been proposed, such as functional analysis of the measurement outcomes, showing a Fisher information increasing linearly with the number of sequential measurements $n_{\mathrm{seq}}$ Burgarth et al. (2015), namely $I_{C}(\lambda)\simeq\frac{n_{\mathrm{seq}}}{\sigma^{2}}\left(\partial_{\lambda}\langle S\rangle_{*}\right)^{2}$, or via correlated stochastic processes Radaelli et al. (2023). Moreover, single-trajectory based sensing Gambetta and Wiseman (2001b); Mabuchi (1996a) utilizing maximum-likelihood estimators Six et al. (2015) (with convergence proof Bompais et al. ) has recently been proposed. This approach demonstrates that a single sequential run (a trajectory) with $n_{\mathrm{seq}}{\gg}1$ is enough to achieve estimation with arbitrary precision. Recently, using a Monte Carlo methodology, a comprehensive analysis of sequential measurements metrology for $n_{\mathrm{seq}}\gg 1$ has been presented. This analysis evaluates the increment in Fisher information asYang et al. (2023):
$\mathcal{F}_{\lambda}^{(n)}{=}\mathcal{F}_{\lambda}^{(n{-}1)}{+}{\Delta}\mathcal{F}_{\lambda}^{(n)};\hskip 14.22636pt\Delta\mathcal{F}_{\lambda}^{(n)}{:=}\sum_{\boldsymbol{\gamma}}P_{\boldsymbol{\gamma}}f_{\lambda}^{\boldsymbol{\gamma},(n)},$ (84)
@@
where $\mathcal{F}_{\lambda}^{(j)}$ represents the CFI at step $j$, $\Delta\mathcal{F}_{\lambda}^{(n)}$ is the increment of the CFI after performing one more measurement following the recording of $(n{-}1)$ measurements, and $f_{\lambda}^{\boldsymbol{\gamma},(n)}$ is the CFI obtained from the $n$-th measurement $p(\gamma_{n})$ in trajectory $\boldsymbol{\gamma}$. Utilizing this approach, one can examine the behavior of the CFI for arbitrarily large numbers of sequential measurements. Conducting local projections on a quantum many-body probe leads to probe’s finite memory, attributed to repeated local projections and free evolution, resulting in a rank-1 matrix Yang et al. (2023). This probe’s finite memory is linked to the transition from super-linear to linear scaling of CFI. After losing memory of an early state, CFI scales linearly with $n_{\mathrm{seq}}$, limiting the extractable information capacity of the quantum probe in sequential measurements.
@@
For the quantum many-body probe described in Eq. (82), Figs. 18(a)-(b) demonstrate a correspondence between the transition from super-linear to linear scaling of CFI and the probe’s finite memory. Specifically, the CFI increment $\Delta F_{B}^{(n_{\mathrm{seq}})}$ becomes nearly constant at around the same $n_{\mathrm{seq}}$ where the averaged fidelity $\langle F\rangle_{\mathrm{traj}}$ also becomes nearly constant. This averaged fidelity measures the similarity between two distinct initial states following the same trajectory Yang et al. (2023) (a single trajectory is composed of $n_{\mathrm{seq}}$ measurement steps). In Fig. 18(c), the CFI $\mathcal{F}_{B}^{(n_{\mathrm{seq}})}$ is plotted as a function of $n_{\mathrm{seq}}$. This plot clearly shows a transition from non-linear to linear behavior, in agreement with the quantum-enhanced sensing shown for short sequential measurements Montenegro et al. (2022) and the linear scaling with indirect methods for large $n_{\mathrm{seq}}$ Burgarth et al. (2014). The Monte Carlo-based methodology discussed above for arbitrarily large $n_{\mathrm{seq}}$ it has been addressed for both quantum many-body probes and light-matter probes Yang et al. (2023).
@@
Figure 18: Sequential measurements metrology for the quantum-many body probe of Eq. (82) and large number of consecutive measurements $n_{\mathrm{seq}}$: $(a)$ CFI increment $\Delta F_{B}^{(n_{\mathrm{seq}})}$ as a function of $n_{\mathrm{seq}}$, $(b)$ Fidelity averaged over $10^{4}$ different trajectories (each trajectory evaluates the fidelity between two distinct initial states following the same trajectory) as a function of $n_{\mathrm{seq}}$, $(c)$ CFI $\mathcal{F}_{B}^{(n_{\mathrm{seq}})}$ as a function of $n_{\mathrm{seq}}$. Figure taken from Yang et al. (2023).
@@
XIV Global Sensing: Parameter Estimation with Minimal Prior Information

### v3

@@ XII Non-Equilibrium Quantum Sensing: Quantum Many-Body Scars and Sensing
are the scar states of Hamiltonian Eq. (80) with $\phi{=}\pm\pi/2$ Mark and Motrunich (2020). In other words, the eigenstates $|\psi(s)\rangle$, defined by $\hat{H}(\phi{=}{\pm}\pi/2)|\psi(s)\rangle{=}\theta(s{-}L/2)|\psi(s)\rangle$, violate the eigenstate thermalization hypothesis as they show a subvolume-law entanglement growth. Rewriting the initial product state $|\psi_{t=0}\rangle{=}\otimes_{i=1}^{L}(|{+1}\rangle{+}|{-1}\rangle){/}\sqrt{2}$ in terms of the elements of the Dicke-scar subspace results in $|\psi_{t{=}0}\rangle{=}\sum_{s=0}^{L}\sqrt{\dfrac{L!}{2^{L}s!(L-s)!}}|\psi(s)\rangle$. The time-evolved state in this subspace is obtained as $|\psi_{t}^{*}\rangle{=}\otimes_{i=1}^{L}(e^{-it^{*}\theta/2}|{+1}\rangle+e^{it^{*}\theta/2}|{-1}\rangle){/}\sqrt{2}$ results in $\delta\theta{=}1/\sqrt{t^{*}Lt_{\mathrm{tot}}}$, despite the presence of strong interaction. Therefore, the initial state which is prepared in the Dicke-scar subspace of Hamiltonian $\hat{H}(\phi{=}\pm\pi/2)$ can result in a vanishing estimation error via its long-lived coherence property $t^{*}{\rightarrow}{\infty}$. More discussion on the advanced role of quantum many-body scar in quantum-enhanced sensing can be found in Refs. Dooley (2021); Dooley et al. (2023); Desaules et al. (2022).
@@ XIII Non-Equilibrium Quantum Sensing: Sequential Measurements Metrology
Measurement represents another distinct feature that sets apart the classical from the quantum world. Indeed, the concept of wave-function collapse due to projective measurements on the quantum system has sparked several discussions within quantum theory itself Machida and Namiki (1980a, b); Furry (1936); Zeh (1970). In quantum many-body probes with partial accessibility, measurements can only be conducted locally on a subsystem. However, in these scenarios, even though the measurement is local, the wave function collapse is global and affects the entire system’s wave function. This intriguing phenomenon has been the subject of intensive studies Busch et al. (1990); Schmidt and Gemmer (2020); Ban (2021); Skinner et al. (2019b); Block et al. (2022a); Benoist et al. (2023); Haapasalo et al. (2016); Benoist et al. (2019); Burgarth et al. (2014); Rybár and Ziman (2015), such as a unique feature for quenching many-body systems Bayat (2017); Pouyandeh et al. (2014); Ma et al. (2018); Bayat et al. (2018), leading to potential new types of phase transitions Choi et al. (2020); Feng et al. (2023); Minato et al. (2022); Skinner et al. (2019a); Gullans and Huse (2020); Block et al. (2022b). These measurement-induced phase transitions have also been recently linked with quantum-enhanced sensitivity (Paviglianiti and Silva, 2023; Di Fresco et al., 2024).
In conventional sensing strategies, data is collected through independent and identical (IID) probability distributions. This means that after measuring the probe, one needs to reset it to its original quantum state for another round of experiments or, equivalently, use identical copies. Note that the formulation of the Cramér-Rao inequality (see Eq. (2)) assumes the resetting of the probe after each measurement or equivalently using $M$ identical probes at once. A simple yet versatile sensing strategy has been devised using sequential measurements on the probe at regular time intervals Burgarth et al. (2015); Guta and Kiukas (2017); Catana et al. (2015); Montenegro et al. (2022); Mabuchi (1996); van Horssen and Guţă (2015). This innovative sequential sensing scheme is of significant importance in quantum metrology for several reasons: (i) it avoids extremely time-consuming resetting times, (ii) it avoids the need for highly correlated or complex initial states, and (iii) available experimental measurements suffice for evidencing quantum-enhanced sensitivity. The sequential measurement sensing scheme, see Fig. 18, is an iterative sensing protocol:
@@
    At time $\tau_{i}$ a positive operator-valued measure (POVM) $\hat{\Upsilon}_{\gamma_{i}}=\hat{\Pi}^{\dagger}_{\gamma_{i}}\hat{\Pi}_{\gamma_{i}}$, where $\hat{\Upsilon}_{\gamma_{i}}\geq 0$ and $\{\hat{\Upsilon}_{\gamma_{i}}\}$ are the elements of the POVM with random outcome $\gamma_{i}$, is performed on the probe, updating the quantum state into
    $\rho^{(i{+}1)}(0){=}\frac{\hat{\Pi}_{\gamma_{i}}\rho^{(i)}(\tau_{i})\hat{\Pi}_{\gamma_{i}}^{\dagger}}{p(\gamma_{i})},$ (83)
@@
    $p(\gamma_{i}){=}\text{Tr}[\hat{\Pi}_{\gamma_{i}}\rho^{(i)}(\tau_{i})\hat{\Pi}_{\gamma_{i}}^{\dagger}],$ (84)
@@
In general, the data collected from sequential measurements are non-IID, meaning that the measurement data are not independent of each other—i.e., one deals with effectively different probes at each measurement step. It is essential to emphasize that, unlike conventional sensing schemes, there is no need for specific maximally entangled probe, the need for supporting quantum phases, feedback mechanisms, or quantum control. Naturally, standard sensing scheme reduces to the particular case of $n_{\mathrm{seq}}{=}1$.
Figure 18: Comparison between conventional and sequential measurements sensing strategies: $(a)$ Conventional strategy for estimating a parameter $g$ of a quantum system, where measurement data $\{s_{1},\ldots,s_{N}\}$ are collected by independent and identical experiments. Every time the experiment is performed, the system is reset to some specific known initial state $\rho_{0}$. $(b)$ Sequential scheme for estimating a parameter $g$ of a quantum system, where the measurements are performed sequentially to collect data $\{s_{1},\ldots,s_{N}\}$ without resetting the state of the system every after the measurement and the initial state $\rho_{0}$ can be arbitrary. Figure taken from Burgarth et al. (2015).
By harnessing the sequential measurements followed by free evolution at regular time intervals, the sequential sensing scheme has led to significant advancements, including Hamiltonian identification Mabuchi (1996), sequential measurements sensing schemes Li and Cui (2023); O’Connor et al. (2024); Clark et al. (2019a); Yang et al. (2024a); De Pasquale et al. (2017a); Clark et al. (2019b); Rasbi et al. (2022); Burgarth et al. (2015); Catana et al. (2015); Guţă (2011); De Pasquale et al. (2017b); Ritboon et al. (2022); Bompais and Amini (2023); Bompais et al. (2022); Gherardini et al. (2019); Gambetta and Wiseman (2001a); Müller et al. (2016); Nagali et al. (2012); Kiilerich and Mølmer (2015).
@@
$\hat{H}=-J\sum_{j=1}^{N-1}\hat{\bm{\sigma}}_{j}\cdot\hat{\bm{\sigma}}_{j+1}+B_{x}\hat{\sigma}_{1}^{x}+B_{z}\hat{\sigma}_{1}^{z},$ (85)
where $\hat{\bm{\sigma}}_{j}{=}(\hat{\sigma}_{j}^{x}{,}\hat{\sigma}_{j}^{y}{,}\hat{\sigma}_{j}^{z})$ represents a vector composed of Pauli matrices acting on qubit at site $j$, $J$ denotes the exchange interaction, and $\bm{B}{=}(B_{x}{,}0{,}B_{z})$ stands for the local magnetic field to be estimated located at first site. For the sake of simplicity, a ferromagnetic state $|\psi(0)\rangle=|\downarrow\downarrow\downarrow\ldots\rangle$ is probed at site $N$ in the $\hat{\sigma}_{z}$ basis at regular time intervals $J\tau=N$. The latter allows enough time for the quantum correlations to spread throughout the spin chain. It has been shown Montenegro et al. (2022) that a synchronization-like magnetization dynamics occurs between the first and last sites. This opens up the possibility of extracting information about the local magnetic field (located at the first site) by observing the dynamics of the readout qubit (located at the last $N$ site). Remote sensing holds special significance in biology, where non-invasive sampling of organic tissues is desirable to protect the biological material (see quantum illumination techniques Albarelli et al. (2020a); Casariego et al. (2022); Reichert et al. (2023) and networks Zhao et al. (2025)).
@@
$t_{\mathrm{tot}}=M(t_{\mathrm{init}}+t_{\mathrm{evo}}+t_{\mathrm{meas}}n_{\mathrm{seq}}),$ (86)
@@
Nonetheless, sequential measurement sensing schemes rapidly become limited to short measurement sequences due to the exponential growth of measurement outcomes with the number of sequences—$(2^{N})^{q}$ for $N$ measurements on $q$ qubits. Thus, evaluating all probabilities distributions becomes intractable. For estimating sensing precision with a large number of sequential measurements, indirect approaches have been proposed, such as functional analysis of the measurement outcomes, showing a Fisher information increasing linearly with the number of sequential measurements $n_{\mathrm{seq}}$ Burgarth et al. (2015), namely $I(\theta)\simeq n_{\mathrm{seq}}$, or via correlated stochastic processes Radaelli et al. (2023). Moreover, single-trajectory based sensing Gambetta and Wiseman (2001b); Mabuchi (1996) utilizing maximum-likelihood estimators Six et al. (2015) (with convergence proof Bompais et al. (2022)) has recently been proposed. This approach demonstrates that a single sequential run (a trajectory) with $n_{\mathrm{seq}}{\gg}1$ is enough to achieve estimation with arbitrary precision. Recently, using a Monte Carlo methodology, a comprehensive analysis of sequential measurements metrology for $n_{\mathrm{seq}}\gg 1$ has been presented. This analysis evaluates the increment in Fisher information as Yang et al. (2023) (for clarity of presentation, we have omitted the unknown parameter to be estimated, that is $I^{(n_{\mathrm{seq}})}(\theta){:=}I^{(n_{\mathrm{seq}})}$):
$I^{(n)}{=}I^{(n{-}1)}{+}{\Delta}I^{(n)};\hskip 14.22636pt\Delta I^{(n)}{:=}\sum_{\boldsymbol{\gamma}}P_{\boldsymbol{\gamma}}f^{\boldsymbol{\gamma},(n)},$ (87)
@@
where $I^{(j)}$ represents the CFI at step $j$, $\Delta I^{(n)}$ is the increment of the CFI after performing one more measurement following the recording of $(n{-}1)$ measurements, and $f^{\boldsymbol{\gamma},(n)}(\theta){:=}f^{\boldsymbol{\gamma},(n)}$ is the CFI obtained from the $n$-th measurement $p(\gamma_{n})$ in trajectory $\boldsymbol{\gamma}$. Utilizing this approach, one can examine the behavior of the CFI for arbitrarily large numbers of sequential measurements. Conducting local projections on a quantum many-body probe leads to probe’s finite memory, attributed to repeated local projections and free evolution, resulting in a rank-1 matrix Yang et al. (2023). This probe’s finite memory is linked to the transition from super-linear to linear scaling of CFI. After losing memory of an early state, CFI scales linearly with $n_{\mathrm{seq}}$, limiting the extractable information capacity of the quantum probe in sequential measurements.
@@
For the quantum many-body probe described in Eq. (85), Figs. 19(a)-(b) demonstrate a correspondence between the transition from super-linear to linear scaling of CFI and the probe’s finite memory. Specifically, the CFI increment $\Delta I^{(n_{\mathrm{seq}})}$ becomes nearly constant at around the same $n_{\mathrm{seq}}$ where the averaged fidelity $\langle F\rangle_{\mathrm{traj}}$ also becomes nearly constant. This averaged fidelity measures the similarity between two distinct initial states following the same trajectory Yang et al. (2023) (a single trajectory is composed of $n_{\mathrm{seq}}$ measurement steps). In Fig. 19(c), the CFI $I^{(n_{\mathrm{seq}})}$ is plotted as a function of $n_{\mathrm{seq}}$. This plot clearly shows a transition from non-linear to linear behavior, in agreement with the quantum-enhanced sensing shown for short sequential measurements Montenegro et al. (2022) and the linear scaling with indirect methods for large $n_{\mathrm{seq}}$ Burgarth et al. (2014). The Monte Carlo-based methodology discussed above for arbitrarily large $n_{\mathrm{seq}}$ it has been addressed for both quantum many-body probes and light-matter probes Yang et al. (2023).
@@
Figure 19: Sequential measurements metrology for the quantum-many body probe of Eq. (85) and large number of consecutive measurements $n_{\mathrm{seq}}$: $(a)$ CFI increment $\Delta I^{(n_{\mathrm{seq}})}$ as a function of $n_{\mathrm{seq}}$, $(b)$ Fidelity averaged over $10^{4}$ different trajectories (each trajectory evaluates the fidelity between two distinct initial states following the same trajectory) as a function of $n_{\mathrm{seq}}$, $(c)$ CFI $I^{(n_{\mathrm{seq}})}$ as a function of $n_{\mathrm{seq}}$. Figure taken from Yang et al. (2023).
In yet another sensing application, sequential measurement sensing directly addresses a fundamental challenge in quantum sensing and metrology: it overcomes the singularity of the QFIM, thereby enabling the simultaneous estimation of multiple parameters Yang et al. (2025). As discussed in Sec. II.4, in multi-parameter quantum sensing, achieving optimal estimation for multiple parameters simultaneously is generally not possible due to the inherent non-commutativity of quantum mechanics Lu and Wang (2021); Carollo et al. (2019a); Di Fresco et al. (2022). As a result, the matrix inequality $\text{Cov}[\tilde{\boldsymbol{\Theta}}]\geq I(\boldsymbol{\theta})^{-1}\geq I_{Q}(\boldsymbol{\theta})^{-1}$ from Eq. (19) is not always saturable Szczykulska et al. (2016); Ragy et al. (2016); Lu and Wang (2021); Albarelli et al. (2020a, 2024); Carollo et al. (2019a); Paris (2009); Demkowicz-Dobrzański et al. (2020); Di Fresco et al. (2022); Belliardo and Giovannetti (2021). Furthermore, the Fisher information matrices often become singular, meaning their inverses do not exist Namkung et al. (2024); Liu et al. (2019); Mihailescu et al. (2024a); Candeloro et al. (2024).
While existing approaches address QFIM singularity using the Moore-Penrose pseudoinverse of the CFIM Namkung et al. (2024) or by formulating the Cramér-Rao bound as an unconstrained quadratic optimization problem Stoica and Marzetta (2001), sequential measurement strategies provide a direct and scalable solution. Indeed, by implementing a sequential measurement protocol with minimal system control Burgarth et al. (2015); Montenegro et al. (2022); Yang et al. (2023); Bompais et al. (2022); Clark et al. (2019b); Ma et al. (2018); Benoist et al. (2019, 2023); Block et al. (2022b); Skinner et al. (2019a), it has been shown that one can always ensure the invertibility of the CFI matrix simply by increasing the length of the measurement sequence Yang et al. (2025). This approach is highly efficient, as the invertibility condition of the CFI matrix improves exponentially with the sequence length.
In summary, as presented throughout this section, sequential measurement metrology addresses several key challenges in quantum sensing and metrology. Notably, (i) it eliminates the need for preparing highly entangled quantum states Burgarth et al. (2015); (ii) allows for system evolution (whether unitary or governed by open quantum dynamics Yang et al. (2024a)) with minimal control Montenegro et al. (2022); (iii) relies solely on local, experimentally feasible measurements Montenegro et al. (2022); and (iv) it enables the simultaneous estimation of multiple parameters by overcoming the QFIM singularity Yang et al. (2025).
One challenge, however, is the practical difficulty of implementing a large number of sequential measurements in an experiment. Yet, as discussed, excessively long measurement sequences would inevitably degrade the encoded information in the quantum probe, making such an approach impractical Yang et al. (2023); Burgarth et al. (2015). Despite the above, experimental progress in sequential measurement techniques has been already pursued. For instance, in 2012, Nagali et al. Nagali et al. (2012) experimentally demonstrated a sequential measurement protocol involving up to two consecutive measurements by exploiting different degrees of freedom of a single photon. Their experimental setup used a Sagnac interferometer with a polarizing beam splitter. Furthermore, using an adaptive strategy, they investigated the optimal trade-off between the maximum extractable information and the disturbance introduced to the system Nagali et al. (2012).
As a final note, sequential measurement sensing may align well with the framework of weak measurements Aharonov et al. (1988); Duck et al. (1989), which has already been shown to enhance metrological capabilities Zhang et al. (2015); Yin et al. (2021); Alves et al. (2015) as well as play a significant role in various other contexts Rozema et al. (2012); Mundarain and Orszag (2016); Feizpour et al. (2011); Lundeen and Steinberg (2009); Coto et al. (2017). In fact, Pfender et al. Pfender et al. (2019) demonstrated that weak measurement-based techniques enable the use of measurement-correlation schemes for detecting extremely weakly coupled single-spin signals. By employing weak measurements, they achieved nuclear magnetic resonance on a 13C nuclear spin at room temperature with a spectral resolution of 3.8 Hz—with the spin undergoing a quantum dynamical phase transition from coherent trapping to coherent oscillation. Similarly, Wang et al. Wang et al. (2019b) characterized correlations in a quantum bath to arbitrary order using a weak measurement protocol. Meinel et al. Meinel et al. (2022), exploiting weak measurements in a nitrogen-vacancy center in diamond, demonstrated that quantum nonlinear spectroscopy can extract arbitrary types and orders of correlations in a quantum system. As shown above, the potential connection between sequential measurement metrology and weak measurements might create experimental opportunities to enhance metrological capabilities.
@@
XIV Global Estimation Theory: Parameter Estimation with Minimal Prior Information